Don operates a taxi business, and this year one of his taxis was damaged in a traffic accident. The taxi was originally purchased for $19500 and the adjusted basis was $1400 at the time of the accident. The taxi was repaired at a cost of $2,800 and insurance reimbursed Don $736 of this cost. What is the amount of Don's casualty loss DEDUCTION

-Bateman Corporation sold an office building that it used in its business for $800,750. Bateman bought the building ten years ago for $599625 and has claimed $201125 of depreciation expense. What is the amount and character of Bateman's gain or loss?
Answer #1

1..

Amount of casualty loss before adjustments

= Lower of adjusted basis at the time of accident or cost repair

= Lower of 1400 or 2800= 1400

Amount of casualty loss deduction = 1400 - reimbursement from insurance 736

= $664

2..

For Corporations, Section 291 recapture is 20% of the lesser of Depreciation taken or the realized gain as ordinary income. The remaining gain is section 1231.

For Bateman Corporation:

Section 291 recapture:

Sale of Office building = $800,750

Less: Purchase of Building = $599,625

Less: Depreciation claimed = $201,125

Realized gain = $402,250 ($800,750 - $398,500)

20% of Realized gain = 0.2 * 402,250= $80,450

20% of depreciation = 0.2 * 201,125= $40,225

Lesser of above two option is $40,225

Hence, Section 291 recapture Ordinary gain = $40,225

Section 1231 gain:

Remaining realized gain = $398,500 - $40,225

Section 1231 gain = $358,275